OpenStudy (kimiroc):

Find the zero function of f(x)=2x+7 Determine the domain of the function g(x)= x+2?x^2-9 A. {x|x≠3} B. {x|x≠0, x≠±3} C. {x|x≠±3} D. {x|x≠0,x≠9}

OpenStudy (swissgirl):

I am assuming that the 0 function is when f(x)=0 which means that 2x+7=0 2x=-7 x=-7/2

OpenStudy (kimiroc):

For the First one I keep getting x=-7/2

OpenStudy (swissgirl):

which is correct. The function in the second part has a ? what does that mean

OpenStudy (kimiroc):

Oh okay so I wasn't wrong lol

OpenStudy (kimiroc):

A?

OpenStudy (swissgirl):

g(x)= x+2?x^2-9 y does it have a question mark?

OpenStudy (swissgirl):

ohhh its \( \huge \frac{x+2} {x^2-9} \)

OpenStudy (kimiroc):

Oh that was was a typo, its suppose to be typed like this....g(x)= x+2/x^2-9

OpenStudy (kimiroc):

Yes, like that you beat me to it lol :)

OpenStudy (swissgirl):

ok so then \(x^2-9=0\) Since if the denomanator equals zero then it goes to infinity and stuff that we dont want \(x^2=9\) \( \sqrt{x^2}=\sqrt{9}\) \(x= \pm 3\)

OpenStudy (swissgirl):

So its option number 3

OpenStudy (swissgirl):

since (3)^2=9 and (-3)^2=9

OpenStudy (kimiroc):

Oh I see so but what's the difference between choice A and C?

OpenStudy (swissgirl):

c is +3 and -3

OpenStudy (swissgirl):

A is only +3

OpenStudy (kimiroc):

Oh nvm I see now.... Thanks a million, you're giving better explanations than my teacher lol!

OpenStudy (swissgirl):

No problem :)